# Uncertainty Quantification for LLM-Based Survey Simulations

### v1

@@ Abstract
We investigate the reliable use of simulated survey responses from large language models (LLMs) through the lens of uncertainty quantification. Our approach converts synthetic data into confidence sets for population parameters of human responses, addressing the distribution shift between the simulated and real populations. A key innovation lies in determining the optimal number of simulated responses: too many produce overly narrow confidence sets with poor coverage, while too few yield excessively loose estimates. To resolve this, our method adaptively selects the simulation sample size, ensuring valid average-case coverage guarantees. It is broadly applicable to any LLM, irrespective of its fidelity, and any procedure for constructing confidence sets. Additionally, the selected sample size quantifies the degree of misalignment between the LLM and the target human population. We illustrate our method on real datasets and LLMs.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in mimicking human behaviors. Recent studies have leveraged LLMs to simulate human responses in various domains, including economic and social science experiments (Aher et al., 2023; Horton, 2023; Chen et al., 2023; Bisbee et al., 2024; Huang et al., 2024; Yang et al., 2024; Ziems et al., 2024), market research (Brand et al., 2023; Gui and Toubia, 2023; Goli and Singh, 2024; Wang et al., 2024), education (Zelikman et al., 2023; Lu and Wang, 2024), and so on. Compared to traditional survey methods that recruit and query real people, LLM simulations offer significant advantages in terms of time and cost efficiency, enabling the generation of large-scale synthetic responses with minimal effort.
@@
Additionally, our work connects to the line of work on model-free statistical inference (Shafer and Vovk, 2008; Bates et al., 2021; Angelopoulos et al., 2023). At a high level, these methods use labeled data from the true distribution to calibrate imperfect point predictions from an arbitrary black-box model and then construct valid set estimates. Our approach follows a similar spirit. The “features” and “labels” in our setting correspond to the survey questions and their population statistics of human responses, respectively. However, our labels are not directly observable. As a result, the labeled calibration data needed for these methods is not available. Moreover, for every simulation sample size $k$, one can produce a point prediction of the label using $k$ synthetic responses generated by the LLM. As the optimal sample size is not known a priori, there are infinitely many candidate point predictions to choose from. This makes it difficult to apply existing methods.
@@
The rest of the paper is organized as follows. Section 2 studies one-dimensional mean estimation as a warm-up example. Section 3 presents the general problem setup and methodology. Section 4 illustrates our proposed method on real datasets. Section 5 concludes the paper.
@@
2 Warm-up: Simulation of Binary Responses
@@ 2.1 Motivating Example: Educational Test
Due to the misalignment between the LLM and students, the distribution of the synthetic data $\{y^{\mathsf{syn}}_{i}\}_{i=1}^{k}$ may be very different from the true response distribution. In this case, the sample mean $\bar{y}^{\mathsf{syn}}$ can be a poor estimate of $\mu$, and $\mathcal{I}^{\mathsf{syn}}(k)$ is generally not a valid confidence interval for $\mu$. In particular, as $k\to\infty$, the interval concentrates tightly around the synthetic mean $\mathbb{E}[y^{\mathsf{syn}}_{1}]$ and fails to cover the true mean $\mu$. When $k$ is small, the interval becomes too wide to be informative, even though it may cover $\mu$ with high probability. We provide an illustration in Figure 1.
Figure 1: The coverage-width trade-off for the simulation sample size $k$. The true distribution is ${\rm Ber}(0.4)$ and the synthetic distribution is ${\rm Ber}(0.6)$. The red dotted horizontal line plots the true mean $\mu=0.4$. The blue line plots the sample mean $\bar{y}^{\mathsf{syn}}_{k}$ of the synthetic data, and the blue shaded region visualizes the confidence interval $\mathcal{I}^{\mathsf{syn}}(k)$, for $k\in[40]$. When $k$ is too small (say $k\leq 6$), the interval $\mathcal{I}^{\mathsf{syn}}(k)$ is too wide. When $k$ is too large (say $k\geq 18$), interval $\mathcal{I}^{\mathsf{syn}}(k)$ fails to cover $\mu$.
@@
Our goal in this work is to develop a principled approach for choosing a good simulation sample size $\widehat{k}$, so that $\mathcal{I}^{\mathsf{syn}}(\widehat{k})$ is a valid confidence interval for $\mu$ while having a modest width. Solving this problem has the following important implications.
@@
    The sample size $\widehat{k}$ also reflects the richness of the human population captured by the LLM. We make an analogy using the classical theory of parametric bootstrap. If a model is trained via maximum likelihood estimation over $k$ i.i.d. human samples, then when performing bootstrap for uncertainty quantification, the bootstrap sample size is usually set to be $k$. In this regard, our approach can be thought of as uncovering the size of human population that the LLM can represent:
    Using the LLM to simulate survey responses is roughly as accurate as surveying $\widehat{k}$ real people.
    We provide an illustration in Figure 2. The larger $\widehat{k}$ is, the more diversity that the LLM appears to capture. In contrast, a small $\widehat{k}$ could imply the peculiarity of the LLM compared to the major population.
@@
Existing works typically measure LLM misalignment using integral probability metrics and $f$-divergences (Santurkar et al., 2023; Dominguez-Olmedo et al., 2024; Durmus et al., 2024), which do not carry operational meanings themselves and at times can be hard to interpret. In contrast, our simulation sample size $\widehat{k}$ is scale-free and has a clear operational meaning.
Figure 2: An interpretation of the simulation sample size $k$ as the size of human population that an LLM can represent. Generating outputs from the LLM can be thought of as “resampling” from $k$ human samples that make up the LLM. The figure is generated by DALL-E (Ramesh et al., 2021), and borrows ideas from the Mechanical Turk, a chess-playing machine from the 18th century with a human player hidden inside.
2.2 Methodology for Selecting the Simulation Sample Size
We now introduce our method for choosing a good simulation sample size $\widehat{k}$. It makes use of similar test questions for which real students’ results are available. If such data is available, we can compare LLM simulations with real students’ results on these questions, and use it to guide the choice of $k$.
Specifically, we assume access to $m$ test questions similar to the question of interest. For example, they can come from previous tests or a question bank. For $j\in[m]$, the $j$-th test question has been tested on $n_{j}$ real students, with test results $\mathcal{D}_{j}=\{y_{j,i}\}_{i=1}^{n_{j}}$. We also simulate LLM responses $\mathcal{D}^{\mathsf{syn}}_{j}=\{y^{\mathsf{syn}}_{j,i}\}_{i=1}^{K}$ to the $j$-th test question, and $\mathcal{D}^{\mathsf{syn}}=\{y^{\mathsf{syn}}_{i}\}_{i=1}^{K}$ to the new test question. Here $K\in\mathbb{Z}_{+}$ is the simulation budget.
@@ 2.2 Methodology for Selecting the Simulation Sample Size
where $\bar{y}^{\mathsf{syn}}_{j,k}=\frac{1}{k}\sum_{i=1}^{k}y^{\mathsf{syn}}_{j,i}$ is the sample mean of the first $k$ samples in $\mathcal{D}^{\mathsf{syn}}_{j}$, and $s^{\mathsf{syn}}_{j,k}=\sqrt{\bar{y}^{\mathsf{syn}}_{j,k}(1-\bar{y}^{\mathsf{syn}}_{j,k})}$ is the estimated standard deviation. We also set the convention $\mathcal{I}^{\mathsf{syn}}(0)=\mathcal{I}^{\mathsf{syn}}_{j}(0)=\mathbb{R}$, as nothing can be said about the true parameter without data. We will pick $\widehat{k}\in\{0,1...,K\}$ such that $\mathcal{I}^{\mathsf{syn}}_{j}(\widehat{k})$ covers $\mu_{j}$ with high probability. We expect this choice of $\widehat{k}$ to be also good for $\mathcal{I}^{\mathsf{syn}}(k)$, as the test questions are similar.
@@ 2.3 Theoretical Analysis
We use $\psi$ and $\{\psi_{j}\}_{j=1}^{m}$ to refer to the test question of interest and the $m$ similar ones, respectively. Students’ performance on test questions are characterized by a performance function $F$: a student with profile $z\in\mathcal{Z}$ answers a question $\psi$ correctly with probability $F(z,\psi)\in[0,1]$. The average student performance on the test questions $\psi$ and $\{\psi_{j}\}_{j=1}^{m}$ are then $\mu=\mathbb{E}_{z\sim\mathcal{P}}F(z,\psi)$ and $\mu_{j}=\mathbb{E}_{z\sim\mathcal{P}}F(z,\psi_{j})$, respectively. In addition, the LLM generates synthetic student performance from a synthetic performance function $F^{\mathsf{syn}}$: when prompted with a synthetic profile $z^{\mathsf{syn}}\in\mathcal{Z}$, the LLM answers a question $\psi$ correctly with probability $F^{\mathsf{syn}}(z^{\mathsf{syn}},\psi)\in[0,1]$.
@@
We are now ready to state the theoretical guarantee of our approach. Its proof is deferred to Section B.2. We note that the assumption $\mathbb{P}(\mu_{j}\geq\bar{y}_{j}\mid\psi_{j})=1/2$ is a CLT approximation and is for mathematical convenience only.
@@
Let Assumptions 2.1 and 2.2 hold. Assume that $\mathbb{P}(\bar{y}_{j}\leq\mu_{j}\mid\psi_{j})=1/2$ for each $j\in[m]$. Fix $\alpha\in(0,1)$. Then the simulation sample size $\widehat{k}$ defined by (2.7) satisfies
$\mathbb{P}\Big{(}\mu\in\mathcal{I}^{\mathsf{syn}}(\widehat{k})\Big{)}\geq 1-\alpha-\sqrt{\frac{2}{m}}.$
The probability is taken with respect to randomness of $\big{\{}(\psi_{j},\mathcal{D}_{j},\mathcal{D}^{\mathsf{syn}}_{j})\big{\}}_{j=1}^{m}$, $\psi$ and $\mathcal{D}^{\mathsf{syn}}$.
On average, the chosen simulation sample size $\widehat{k}$ leads to a confidence interval $\mathcal{I}^{\mathsf{syn}}(\widehat{k})$ that covers the true mean $\mu$ with probability at least $1-\alpha-O(\sqrt{1/m})$. As $m\to\infty$, the aforementioned lower bound converges to $1-\alpha$.
@@ 2.4 Sharpness of Sample Size Selection
Consider the setting in Section 2.3 with the following modifications. On a test question $\psi\in\Psi$, the performance (e.g., score) of a real student with profile $z$ follows the distribution $N(F(z,\psi),1)$ instead of $\mathrm{Bernoulli}(F(z,\psi))$; similarly, the performance of the LLM prompted with synthetic profile $z^{\mathsf{syn}}$ has distribution $N(F^{\mathsf{syn}}(z^{\mathsf{syn}},\psi),1)$. Moreover, the confidence intervals $\mathcal{I}^{\mathsf{syn}}(k)$ defined in (2.3) and $\mathcal{I}^{\mathsf{syn}}_{j}(k)$ defined in (2.4) are changed to
$\mathcal{I}^{\mathsf{syn}}(k)=\left[\bar{y}^{\mathsf{syn}}_{k}-\frac{C}{\sqrt{k}},~{}\bar{y}^{\mathsf{syn}}_{k}+\frac{C}{\sqrt{k}}\right],$
$\mathcal{I}^{\mathsf{syn}}_{j}(k)=\left[\bar{y}^{\mathsf{syn}}_{j,k}-\frac{C}{\sqrt{k}},~{}\bar{y}^{\mathsf{syn}}_{j,k}+\frac{C}{\sqrt{k}}\right],$
@@
$\widehat{k}\geq\min\left\{\left(\frac{C}{5\Delta}\right)^{2},n,K\right\}.$
When this happens, the selected confidence interval $\mathcal{I}^{\mathsf{syn}}(\widehat{k})$ has width $O\left(\max\{\Delta,n^{-1/2},K^{-1/2}\}\right)$.
Theorem 2.2 implies that the chosen $\widehat{k}$ is the largest possible, or equivalently, the interval $\mathcal{I}^{\mathsf{syn}}(\widehat{k})$ is the shortest possible. To see this, suppose that the simulation budget $K$ is large. When $\Delta$ is small, there is little distribution shift between the real and simulated student performance, so it is optimal to use as many synthetic samples as possible. Theorem 2.2 shows that in this case, the procedure (2.7) chooses $\widehat{k}=n$ with high probability. This is the maximum effective sample size, since the $n$ real data points can identify the true mean only up to an error of $O(n^{-1/2})$ and thus there is no statistical benefit in simulating more than $n$ samples. On the other hand, when the discrepancy $\Delta$ between the real and simulated student performance is large, any $\mathcal{I}^{\mathsf{syn}}(k)$ that covers the true mean with high probability must have width $\Omega(\Delta)$. Thus, the chosen confidence interval $\mathcal{I}^{\mathsf{syn}}(\widehat{k})$, which has width $O(\Delta)$ with high probability, is the shortest possible.

### v3

@@ Abstract
We investigate the use of large language models (LLMs) to simulate human responses to survey questions, and perform uncertainty quantification to gain reliable insights. Our approach converts imperfect LLM-simulated responses into confidence sets for population parameters of human responses, addressing the distribution shift between the simulated and real populations. A key innovation lies in determining the optimal number of simulated responses: too many produce overly narrow confidence sets with poor coverage, while too few yield excessively loose estimates. To resolve this, our method adaptively selects the simulation sample size, ensuring valid average-case coverage guarantees. It is broadly applicable to any LLM, irrespective of its fidelity, and any procedure for constructing confidence sets. Additionally, the selected sample size quantifies the degree of misalignment between the LLM and the target human population. We illustrate our method on real datasets and LLMs.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in mimicking human behaviors. Recent studies have leveraged LLMs to simulate human responses in various domains, including economic and social science experiments (Aher et al., 2023; Horton, 2023; Chen et al., 2023; Bisbee et al., 2024; Huang et al., 2024; Yang et al., 2024; Ziems et al., 2024), market research (Brand et al., 2023; Gui and Toubia, 2023; Goli and Singh, 2024; Wang et al., 2024), education (Zelikman et al., 2023; Lu and Wang, 2024), and so on. The typical simulation procedure consists in prompting an LLM with a real or fictional persona as well as a survey question, and collecting the LLM’s responses. Compared to traditional survey methods that recruit and query real people, LLM simulations offer significant advantages in terms of time and cost efficiency, enabling the generation of large-scale synthetic responses with minimal effort.
@@
Additionally, our work connects to the line of work on model-free statistical inference, including conformal inference (Vovk et al., 2005; Shafer and Vovk, 2008; Bates et al., 2021; Angelopoulos et al., 2024; Kim et al., 2024) and prediction-powered inference (Angelopoulos et al., 2023). At a high level, these methods use labeled data from the true distribution to calibrate imperfect point predictions from an arbitrary black-box model and then construct valid set estimates. Our approach follows a similar spirit. The “features” and “labels” in our setting correspond to the survey questions and their population statistics of human responses, respectively. However, our labels are not directly observable. As a result, the labeled calibration data needed for these methods is not available. Moreover, for every simulation sample size $k$, one can produce a point prediction of the label using $k$ synthetic responses generated by the LLM. As the optimal sample size is not known a priori, there are infinitely many candidate point predictions to choose from. This makes it difficult to apply existing statistical inference methods.
@@
The rest of the paper is organized as follows. Section 2 studies binary response simulation as a warm-up example. Section 3 presents the general problem setup and methodology. Section 4 illustrates our proposed method on real datasets. Section 5 concludes the paper.
@@
2 Warm-up: Binary Response Simulation
@@ 2.1 Motivating Example: Educational Test
Due to the misalignment between the LLM and students, the distribution of the synthetic data $\{y^{\mathsf{syn}}_{i}\}_{i=1}^{k}$ may be very different from the true response distribution. In this case, the sample mean $\bar{y}^{\mathsf{syn}}$ can be a poor estimate of $\mu$, and $\mathcal{I}^{\mathsf{syn}}(k)$ is generally not a valid confidence interval for $\mu$. In particular, as $k\to\infty$, the interval concentrates tightly around the synthetic mean $\mathbb{E}[y^{\mathsf{syn}}_{1}]$ and fails to cover the true mean $\mu$. On the other hand, when $k$ is small, the interval becomes too wide to be informative, even though it may cover $\mu$ with high probability. We provide an illustration in Figure 1.
Figure 1: The coverage-width trade-off for the simulation sample size $k$. The true distribution is $\mathrm{Bernoulli}(0.4)$ and the synthetic distribution is $\mathrm{Bernoulli}(0.6)$. The horizontal axis is the simulation sample size $k$. The red dotted horizontal line plots the true mean $\mu=0.4$. The blue curve plots the sample mean $\bar{y}^{\mathsf{syn}}_{k}$ of the synthetic data, and the blue shaded region visualizes the confidence interval $\mathcal{I}^{\mathsf{syn}}(k)$, for $k\in[40]$. When $k$ is too small (say $k\leq 6$), the interval $\mathcal{I}^{\mathsf{syn}}(k)$ is too wide. When $k$ is too large (say $k\geq 18$), $\mathcal{I}^{\mathsf{syn}}(k)$ fails to cover $\mu$.
@@
Our goal in this work is to develop a principled approach for choosing a good simulation sample size $\widehat{k}$, so that $\mathcal{I}^{\mathsf{syn}}(\widehat{k})$ is a valid confidence interval for $\mu$ while having a small width. Solving this problem has the following important implications.
@@
    The sample size $\widehat{k}$ reflects the size of the target population that the LLM can represent. We make an analogy using the classical theory of parametric bootstrap. Suppose a model is trained via maximum likelihood estimation over $k$ i.i.d. human samples. When performing parametric bootstrap for uncertainty quantification, the bootstrap sample size is usually set to be the training sample size $k$. Thus, our simulation sample size $\widehat{k}$ reveals the LLM as “being made up of” $\widehat{k}$ people from the population. We provide a visualization in Figure 4 of Appendix A. The larger $\widehat{k}$ is, the more diversity that the LLM appears to capture. In contrast, a small $\widehat{k}$ could imply the peculiarity of the LLM compared to the major population.
@@
Existing works typically measure LLM misalignment using integral probability metrics and $f$-divergences (Santurkar et al., 2023; Dominguez-Olmedo et al., 2024; Durmus et al., 2024), which do not carry operational meanings themselves and can be hard to interpret. In contrast, our simulation sample size $\widehat{k}$ provides actionable guidance and is easy to understand.
2.2 Methodology for Selecting Simulation Sample Size
We now introduce our method for choosing a good simulation sample size $\widehat{k}$. It makes use of similar test questions for which real students’ results are available. If such data is available, then we can compare LLM simulations with real students’ results on these questions, and use it to guide the choice of $k$.
Specifically, we assume access to $m$ test questions similar to the question of interest. For example, they can come from previous tests or a question bank. For $j\in[m]$, the $j$-th test question has been tested on $n_{j}$ real students, with test results $\mathcal{D}_{j}=\{y_{j,i}\}_{i=1}^{n_{j}}$. We also simulate LLM responses $\mathcal{D}^{\mathsf{syn}}_{j}=\{y^{\mathsf{syn}}_{j,i}\}_{i=1}^{K}$ to the $j$-th test question, and $\mathcal{D}^{\mathsf{syn}}=\{y^{\mathsf{syn}}_{i}\}_{i=1}^{K}$ to the test question of interest. Here $K\in\mathbb{Z}_{+}$ is the simulation budget.
@@ 2.2 Methodology for Selecting Simulation Sample Size
where $\bar{y}^{\mathsf{syn}}_{j,k}=\frac{1}{k}\sum_{i=1}^{k}y^{\mathsf{syn}}_{j,i}$ is the sample mean of the first $k$ samples in $\mathcal{D}^{\mathsf{syn}}_{j}$, and $s^{\mathsf{syn}}_{j,k}=\sqrt{\bar{y}^{\mathsf{syn}}_{j,k}(1-\bar{y}^{\mathsf{syn}}_{j,k})}$ is the estimated standard deviation. We also set the convention $\mathcal{I}^{\mathsf{syn}}(0)=\mathcal{I}^{\mathsf{syn}}_{j}(0)=[0,1]$, as nothing can be said about the true parameter without data. We will pick $\widehat{k}\in\{0,1...,K\}$ such that $\mathcal{I}^{\mathsf{syn}}_{j}(\widehat{k})$ covers $\mu_{j}$ with high probability. We expect this choice of $\widehat{k}$ to be also good for $\mathcal{I}^{\mathsf{syn}}(k)$, as the test questions are similar.
@@ 2.3 Theoretical Analysis
We use $\psi$ and $\{\psi_{j}\}_{j=1}^{m}$ to denote the test question of interest and the $m$ similar ones, respectively. Students’ performance on test questions is characterized by a performance function $F$: a student with profile $z\in\mathcal{Z}$ answers a question $\psi$ correctly with probability $F(z,\psi)\in[0,1]$. The average student performance on the test questions $\psi$ and $\{\psi_{j}\}_{j=1}^{m}$ are then $\mu=\mathbb{E}_{z\sim\mathcal{P}}F(z,\psi)$ and $\mu_{j}=\mathbb{E}_{z\sim\mathcal{P}}F(z,\psi_{j})$, respectively. In addition, the LLM generates synthetic student performance from a synthetic performance function $F^{\mathsf{syn}}$: when prompted with a synthetic profile $z^{\mathsf{syn}}\in\mathcal{Z}$, the LLM answers a question $\psi$ correctly with probability $F^{\mathsf{syn}}(z^{\mathsf{syn}},\psi)\in[0,1]$.
@@
We are now ready to state the theoretical guarantee of our approach. Its proof is deferred to Section B.2. The assumption $\mathbb{P}(\bar{y}_{j}\leq\mu_{j}\mid\psi_{j})\in[\frac{1}{2}-\eta,\frac{1}{2}+\eta]$ quantifies the CLT approximation $\mathbb{P}(\bar{y}_{j}\leq\mu_{j}\mid\psi_{j})\approx 1/2$.
@@
Let Assumptions 2.1 and 2.2 hold. Assume that $\mathbb{P}(\bar{y}_{j}\leq\mu_{j}\mid\psi_{j})\in[\frac{1}{2}-\eta,\frac{1}{2}+\eta]$ for each $j\in[m]$, where $\eta\in[0,1/2)$. Fix $\alpha\in(0,1)$. Then the simulation sample size $\widehat{k}$ defined by (2.7) satisfies
$\mathbb{P}\Big{(}\mu\in\mathcal{I}^{\mathsf{syn}}(\widehat{k})\Big{)}\geq 1-(1-2\eta)^{-1}\bigg{(}\alpha+\sqrt{\frac{2}{m}}\bigg{)}.$
The probability is taken with respect to the randomness of $\big{\{}(\psi_{j},\mathcal{D}_{j},\mathcal{D}^{\mathsf{syn}}_{j})\big{\}}_{j=1}^{m}$, $\psi$ and $\mathcal{D}^{\mathsf{syn}}$.
To interpret Theorem 2.1, suppose $\eta=0$ for simplicity. Then on average, the chosen simulation sample size $\widehat{k}$ leads to a confidence interval $\mathcal{I}^{\mathsf{syn}}(\widehat{k})$ that covers the true mean $\mu$ with probability at least $1-\alpha-O(\sqrt{1/m})$. As $m\to\infty$, the aforementioned lower bound converges to $1-\alpha$.
@@ 2.4 Sharpness of Sample Size Selection
Consider the setting in Section 2.3 with the following modifications. On a test question $\psi\in\Psi$, the performance (e.g., score) of a real student follows a Gaussian distribution with mean $\mathbb{E}_{z\sim\mathcal{P}}F(z,\psi)$ and variance $1$, instead of a Bernoulli distribution with mean $\mathbb{E}_{z\sim\mathcal{P}}F(z,\psi)$. Similarly, the performance of the LLM follows a Gaussian distribution with mean $\mathbb{E}_{z^{\mathsf{syn}}\sim\mathcal{P}^{\mathsf{syn}}}F^{\mathsf{syn}}(z^{\mathsf{syn}},\psi)$ and variance $1$. Moreover, the confidence intervals $\mathcal{I}^{\mathsf{syn}}(k)$ defined in (2.3) and $\mathcal{I}^{\mathsf{syn}}_{j}(k)$ defined in (2.4) are changed to
$\mathcal{I}^{\mathsf{syn}}(k)=\left[\bar{y}^{\mathsf{syn}}_{k}-\frac{C}{\sqrt{k}},~{}\bar{y}^{\mathsf{syn}}_{k}+\frac{C}{\sqrt{k}}\right]\quad\text{and}\quad\mathcal{I}^{\mathsf{syn}}_{j}(k)=\left[\bar{y}^{\mathsf{syn}}_{j,k}-\frac{C}{\sqrt{k}},~{}\bar{y}^{\mathsf{syn}}_{j,k}+\frac{C}{\sqrt{k}}\right],$
@@
$\widehat{k}\geq\min\left\{K,\,n,\,\left(\frac{C}{5\Delta}\right)^{2}\right\}.$
When this happens, the selected confidence interval $\mathcal{I}^{\mathsf{syn}}(\widehat{k})$ has width $O\left(\max\{K^{-1/2},n^{-1/2},\Delta\}\right)$.
Theorem 2.2 implies that the interval $\mathcal{I}^{\mathsf{syn}}(\widehat{k})$ is the shortest possible. To see this, suppose that the simulation budget $K$ is large, then with high probability, $\mathcal{I}^{\mathsf{syn}}(\widehat{k})$ has width $O(\max\{\Delta,n^{-1/2}\})$. This is the optimal width because of the following reasons. First, in the worst case, any $\mathcal{I}^{\mathsf{syn}}(k)$ that covers the true mean with high probability must have width $\Omega(\Delta)$ in order to address the distribution shift between the real and simulated responses. Second, as $n$ real human responses can identify the true mean up to an error of $O(n^{-1/2})$, then any valid $\mathcal{I}^{\mathsf{syn}}(k)$ must also have width $\Omega(n^{-1/2})$. This shows the sharpness of the chosen sample size $\widehat{k}$ and the confidence interval $\mathcal{I}^{\mathsf{syn}}(\widehat{k})$.